silvernight269:

If x-2 is a factor of x^2-bx+b, where b is a constant, what is the value of b?

jhonyy9:

given x-2 is a factor of x^2 -bx +b so what mean this ?

silvernight269:

Wdym

jhonyy9:

so this mean that you can divide x^2 -bx +b by x-2 do you can make it ?

jhonyy9:

x^2 -bx +b / x-2 = x +(-b+2) x^2 -2x


-- -bx-(-2x) +b -bx +2x +b (-b+2)x +b (-b+2)x -2(-b+2)

b -(-2(-b+2)) b -(2b-4) b-2b+4 -b+4 so from this result that x^2 -bx +b is divisibly by x-2 than -b+4=0 -b+4=0 b=4 do you understand now the way ?

jhonyy9:

@ThisGirlPretty

jhonyy9:

how you think this understandably sure ?

jhonyy9:

ask me what you dont understood it

sillybilly123:

If \(x-2\) is a factor of \(f(x) = x^2-bx+b\), then: \(f(x) = x^2-bx+b = (x - 2) ( x + C)\) ... where \(C\) is some [other] number \(\implies f(2) = (2 - 2) ( 2 + C) = 0 = 2^2 - b(2) + b \) \( \implies b = 4\)